OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is parallel to the given line. 9. (-1,3); y=2x+2 10. (1,2)/ y= -5x+4 11. (18,2); 3y-x= -12 12. (2, -5); 2y=3x+10

OpenStudy (anonymous):

The / on number 10 is supposed to be ;

OpenStudy (aripotta):

parallel lines have the same slope. since we're given a point for each question, we can use the point-slope form: y - y2 = m(x - x2) where m is the slope

OpenStudy (aripotta):

so i'll do 9 for you, and then you can do 10. y = 2x + 2 has a slope of 2. we're given the point (-1,3) y - y2 = m(x - x2) y - 3 = 2(x + 1) that's the answer in point-slope form. if we have to have the answer in slope-intercept form, it's easy to convert. y - 3 = 2(x + 1) y - 3 = 2x + 2 y = 2x + 5

OpenStudy (aripotta):

@Ariaroseb you there?